OpenStudy (anonymous):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 4, -8, and 2 + 5i

OpenStudy (asnaseer):

use the fact that an n'th degree polynomial with roots \(x=r_1,x=r_2,...,x=r_n\) can be written as:\[f(x)=C(x−r_1)(x−r_2)...(x−r_n)\]where C is some constant. in addition, if a polynomial has any complex roots, then they ALWAYS appear in pairs where each pair contains a complex number and its complex conjugate. example: if a polynomial has roots 4 and 1+2i, then we can write it as:\[f(x)=C(x−4)(x−(1+2i))(x-(1-2i))\]since (1-2i) is the complex conjugate of (1+2i)

OpenStudy (anonymous):

would this be right f(x) = x4 - 6x3 + 20x2 - 122x + 928

OpenStudy (helder_edwin):

no, it is not. Following what @asnaseer told you, you should multiply \[ \large (x-4)(x+8)(x-(2+5i))(x-(2-5i)) \]

OpenStudy (anonymous):

f(x) = x4 - 19x2 + 244x - 928

OpenStudy (helder_edwin):

yes, well done

OpenStudy (dumbcow):

a tip for complex roots: first write it as x = 2 +- 5i then get the imaginary part alone x-2 = +-5i square both sides (x-2)^2 = 25i^2 simplify x^2 -4x +4 = -25 set equal to zero x^2 -4x +29 = 0 --> (x-4)(x+8)(x^2-4x+29)

OpenStudy (asnaseer):

nice tip @dumbcow :)

OpenStudy (asnaseer):

and well done @babydoll332 :)

OpenStudy (anonymous):

thank you

OpenStudy (asnaseer):

yw :)

OpenStudy (anonymous):

thank you guys as well :D